OpenStudy (anonymous):

3+6x4 divide by 4 ??

OpenStudy (anonymous):

3/4 + 6 = 6.75

OpenStudy (anonymous):

6*4=24+3+27/4+6.75 ans, is 6.75

OpenStudy (anonymous):